If it takes 50 J of energy to move 10 C of charge from point A to point B, what is the magnitude of the potential difference between points A and B?




1
Expert's answer
2020-09-21T06:23:51-0400

By definition, potential difference "V" between points A and B is work required to move a charge from point A to point B per unit charge.

"\displaystyle V = \frac{W}{q} = \frac{50}{10} = 5" V (5 volt).

Answer: 5 V.